# Spin-dependent Off-Axis Holography in Magnetic Environments

### v1

@@ II Experimental Methods
During the last 25 seconds of cooling, spin-dependent off-axis holographic imaging [19] is performed using a linearly polarized probe beam detuned by $\delta=-175\ \rm{MHz}$ (approximately 18 linewidths) from the $|F=1\rangle$ to $|F^{\prime}=1\rangle$ resonance. In total, 50 images are taken of the atoms at a rate of 2 Hz, with probe pulses 25 $\mu\rm{s}$ long with an intensity of roughly 80 $\mu\rm{W}/\rm{cm}^{2}$, corresponding to a saturation parameter of $s\approx 1.0\times 10^{-5}$. This ensures that the imaging is done well within the linear dispersive regime and does not cause significant additional heating. The final stage of cooling reduces the temperature from 20 $\mu\rm{K}$ to below 1 $\mu\rm{K}$ with the onset of condensation after 20 seconds at a temperature of $\sim 1.03\ \mu\rm{K}$. Following the Fourier-cut analysis described in Ref. [19], a pair of images corresponding to opposite polarizations ($\sigma^{+}$, $\sigma^{-}$) are reconstructed. Examples of such images are shown in Figure 1a, b.
@@ III.1 Polarizability
Note that the reduction of $\overleftrightarrow{\alpha}$ to a diagonal matrix greatly simplifies the analysis. For sodium atoms in the $F=1$ hyperfine ground state, the transition strengths for $\sigma^{+}$ light are proportional to the squares of the Clebsch-Gordan coefficients. These yield total coupling strengths of $5/6$, $4/6$, and $3/6$ for $m=-1,0,+1$, respectively [27], as shown in Fig. 2.
Figure 2: Coupling strengths of the various transitions from the $F=1$ ground state of sodium with right-handed circular polarized light. All strengths have to be normalized to the strongest transitions, which has a strength of 60. Note that the total coupling from the different magnetic substates to all excited states are $\nicefrac{{5}}{{6}}$, $\nicefrac{{4}}{{6}}$, and $\nicefrac{{3}}{{6}}$ for $m=$1, 0, and +1, respectively. Values taken from [27]
To account for the local magnetic field orientation, the ground state density matrix is rotated from the field-aligned frame (where atoms occupy the $m=-1$ state) to the probe beam frame (aligned with $y$). The rotation is
@@ III.2 Magnetic Field Geometry
$K_{ab}^{\prime\prime}$ $-1.38\times 10^{-4}$ G/(A mm2)
@@
$K_{ab}$ $-1.02$ G/A
Figure 3: Orientation of the magnetic field characterized by the angle $\beta$ relative to the $y$-direction in the $xz$-plane, at $y=0$ (color). The dashed ellipse represents the spatial extent of a thermal atomic cloud at $T=30\ \mu\rm{K}$ in in-situ conditions. Here, $\sigma_{z}=\sqrt{k_{B}T/m\omega_{z}^{2}}\approx 1.1$ mm and $\sigma_{r}=\sqrt{k_{B}T/m\omega_{\rho}^{2}}\approx 0.12$ mm, with $m$ in this case denoting the mass of sodium atoms, and $\omega_{z}$, $\omega_{\rho}$ the trapping frequency in axial and radial direction, respectively. The arrows in the right-hand plot indicate the direction of the quantization axis relative to the $y$-direction - the direction of the propagation of light - along the center of the trap ($y=0,\ z=0$).
@@ III.3 Rotation Matrices
where $B_{\mathrm{tot}}=\sqrt{B_{x}^{2}+B_{y}^{2}+B_{z}^{2}}$. To rotate the field vector on the $y$ axis, a rotation by $-\phi$ about $\hat{\mathbf{y}}$ is applied followed by a rotation by $-\theta$ about $\hat{\mathbf{z}}$. The combined transformation in Euler‐angle convention [^2] Euler angles $(\alpha,\beta,\gamma)$ define rotations in the Wigner $D$‐matrix formalism. Note that $\beta$ above (deviation from the $y$ axis) differs from the Euler angle $\beta$. is
@@
with the upper-left element representing $m^{\prime}=-1$, $m=-1$. Applying this transformation to the density matrix yields
@@
where the population is initially in the $m=-1$ state. Equation (20) then becomes:
@@
which satisfy $\sum_{m}\Pi_{m}=1$. For $\theta=0$, the entire population occupies $m=-1$; for $\theta=\pi$, $m=+1$; and for $\theta=\pi/2$, one finds $\Pi_{\pm 1}=1/4$ and $\Pi_{0}=1/2$.
@@ III.4 Accumulated Phase and Column Densities
Figure 4: Simulated phase delay contours for a) $\sigma_{+}$ and b) $\sigma_{-}$ polarizations showing an apparent displacement of the phase centroid in a thermal atomic cloud of $3\times 10^{8}$ atoms at $30\ \rm{\mu K}$. The dashed white line indicates its Gaussian widths. c) Vertical and d) horizontal line cuts through the center of the clouds highlight an apparent radial shift of about ±$\pm 40\ \mu\text{m}$ for $\sigma^{\pm}$-polarized light, relative to the center of the ground truth (GT) phase delay for $\beta=\pi$. Note the different x-axis scales in the plots in c) and d) due to the cigar-shaped geometry of the cloud. The vertical cut is scaled for better visualization of the displacement.
@@
with radial and axial widths $\sigma_{r},\sigma_{z}$, the spatial dependence of $\Pi_{m}(\beta)$ induces an asymmetric phase profile for circularly polarized light. Specifically, for right- versus left-handed circular polarization, the phase contour is displaced oppositely along the radial axis. The simulation in Fig. 4 illustrates this effect for a cloud of $3\times 10^{8}$ atoms at $30\,\mu\text{K}$, exhibiting a radial shift of approximately ±$\pm 40\ \mu\text{m}$ relative to the centroid of the phase distribution at $\beta=\pi$. Although an axial shift is, in principle, also expected, it depends on the magnetic-field gradient along the axial direction, which is negligible near the trap center. As a result, no significant axial displacement is observed in the simulations or in the experiment.
@@ IV Comparison to the Experimental Data
Figure 5: Analysis of the relative shift imposed by the inhomogeneous magnetic field onto two differently polarized circular polarizations of light. a) Relative displacements of the centers of the fitted density distributions show a decreasing pattern, as a function of cooling time. At each stage, the temperature is calculated using fits to the data and a simulation using the model from above is used to compute the displacement. The temperatures at each point are shown in b). The shift, as a function of temperature is shown in c). At high temperatures, the model seems to predict the behavior well, but it seems not to fully account for the shift at lower temperatures. The solid black dots show the data; empty dots with a dashed red guide show the simulation.
@@
×$\times\Bigl{(}1-\tfrac{1}{2}\sin^{2}\beta(x,y,z)\Bigr{)},$ (29)
@@ V Conclusion
We have introduced a compact three–dimensional framework that links the local tilt of the magnetic field to the phase accumulated by the circular components of an off-axis probe. By combining tensor polarizability with field–frame rotations, the model predicts both phase and spin density transformations, provided $\mathbf{B}(x,y,z)$ is known. The theoretical predictions follow experimental SOAH data down to a few $\mu K$, and reveal a systematic displacement of roughly $\sim 1\,\mu\text{m}$ that persists even in a uniform dipole trap. This residual shift has been observed and discussed before [19], and is likely to originate from optical asymmetries rather than magnetic curvature.
Because atomic birefringence acts as a virtual lens, every imaging scheme that relies on computing the difference between $\sigma^{+}$ and $\sigma^{-}$ pixels must correct the resulting aberrations, which are often of the order of tens of $\mu\rm{m}$. The shift can be corrected in post-processing by multiplying the polarization images in the Fourier space with a differential phase mask which can be evaluated as a function of the magnetic field environment. For small clouds and for $\beta\ll 1$, this factor can be seen as roughly linear, allowing for a fast correction even when the magnetic fields are not precisely known. When the magnetic field is known exactly, this method allows for higher-order aberration corrections down to the diffraction limit.

### v2

# Polarization-Sensitive Imaging in Magnetic Environments

@@ II Experimental Methods
To experimentally probe the effects of magnetic fields, we trap and cool sodium atoms in a magnetic trap and monitor them during the final stage of evaporative cooling. Using spin-dependent off-axis holography (SOAH), we image both the density and spin distribution of the cloud through the last 25 seconds of cooling, to and past the Bose–Einstein condensation threshold. Because SOAH records the full optical field, we can choose the probe detuning such that absorptive effects that lead to heating are suppressed by a factor of $\sim 1300$, while the phase signal remains appreciable. This makes it possible to follow a single cloud throughout the experiment without significant particle loss, capturing its evolution across the transition and revealing systematic shifts between the spin-resolved images.
@@
During the last 25 seconds of cooling, spin-resolved imaging is performed by employ spin-dependent off-axis holography (SOAH) [19]. In this method the linearly polarized probe beam interacts with the atomic cloud, such that the two circular components acquire different phase shifts depending on the local spin state. The probe beam is then interfered on the camera with two circularly polarized reference beams. Each of the two reference beams is incident at a different, small angle, which produces spatial interference fringes for both $\sigma^{+}$ and $\sigma^{-}$ beams on the camera in a single exposure. As the two reference beams hit the camera under different angles, the hologram can be Fourier filtered in such a way, to reconstruct the entire complex optical field for each polarization component separately.
In our implementation, the probe beam is detuned by $\delta=-175\ \mathrm{MHz}$ (about $\delta/\Gamma\approx$18 linewidths) from the $|F=1\rangle\rightarrow|F^{\prime}=1\rangle$ resonance. As the absorption scales approximately with $(\Gamma/2\delta)^{2}$, while the phase delay scales with $\Gamma/2\delta$, the relative strength of phase delay compared to absorption is enhanced by a factor of $2\delta/\Gamma\approx 36$. Thus, the effective OD is strongly suppressed, making the imaging minimally destructive, while maintaining good sensitivity to spin-dependent phase shifts.
In total, 50 images are taken of the atoms at a rate of 2 Hz, with probe pulses 25 $\mu\rm{s}$ long with an intensity of roughly 80 $\mu\rm{W}/\rm{cm}^{2}$, corresponding to a saturation parameter of $s\approx 1.0\times 10^{-5}$. This ensures that the imaging is done well within the linear dispersive regime and does not cause significant additional heating. The final stage of cooling reduces the temperature from 20 $\mu\rm{K}$ to below 1 $\mu\rm{K}$ with the onset of condensation after 20 seconds at a temperature of $\sim 1.03\ \mu\rm{K}$. Following the Fourier-cut analysis described in Ref. [19], a pair of images corresponding to opposite polarizations ($\sigma^{+}$, $\sigma^{-}$) are reconstructed. Examples of such images are shown in Figure 1a, b.
@@ III Theoretical Framework
The description begins with the response of the atomic ensemble to an applied electromagentic field. This response is captured by the atomic polarizability, which links the local electric field to the induced atomic polarization. The atomic polarizability ultimately determines both the absorption and phase shift experienced by the probe. In spin-polarized gases such as ours, the response is anisotropic and depends on the internal spin state as well as the orientation of the local magnetic field. In this section, we develop a description of this polarizability, beginning from the general susceptibility tensor and reducing it to the case where a linearly polarized probe decomposes into circular components that couple selectively to $\sigma^{+}$ or $\sigma^{-}$ transitions.
@@ III.1 Polarizability
Note that the reduction of $\overleftrightarrow{\alpha}$ to a diagonal matrix greatly simplifies the analysis. For sodium atoms in the $F=1$ hyperfine ground state, the transition strengths for $\sigma^{+}$ light are proportional to the squares of the Clebsch-Gordan coefficients. These yield total coupling strengths of $5/6$, $4/6$, and $3/6$ for m=1,0,+1m=-1,0,+1, respectively [27], as shown in Fig. 2.
Figure 2: Coupling strengths of the various transitions from the $F=1$ ground state of sodium with right-handed circular polarized light. All strengths have to be normalized to the strongest transitions, which has a strength of 60. Note that the total coupling from the different magnetic substates to all excited states are $\nicefrac{{5}}{{6}}$, $\nicefrac{{4}}{{6}}$, and $\nicefrac{{3}}{{6}}$ for $m=$ -1, 0, and +1, respectively. Values taken from [27]
To account for the local magnetic field orientation, the ground state density matrix is rotated from the field-aligned frame (where atoms occupy the m=1m=-1 state) to the probe beam frame (aligned with $y$). The rotation is
@@ III.2 Magnetic Field Geometry
$K_{ab}^{\prime\prime}$ $1.38\times 10^{-4}$ G/(A mm2)
@@
$K_{ab}$ 1.02-1.02 G/A
In the radial direction, the clover-leaf trap potential transitions from harmonic to linear when the thermal energy equals the Zeeman energy, i.e. when $k_{B}T=\mu B_{0}$. For our operating conditions, with $B_{0}\approx 2~\text{G}$, this crossover occurs at $T\approx 70~\mu\text{K}$. Since the highest temperatures reached in our experiment are on the order of $20~\mu\text{K}$, the radial confinement is already well within the harmonic regime. Along the axial direction, the trap remains harmonic at all temperatures of interest.
Figure 3: Orientation of the magnetic field characterized by the angle $\beta$ relative to the $y$-direction in the xzxz-plane, at $y=0$ (color). The dashed ellipse represents the spatial extent of a thermal atomic cloud at $T=30\ \mu\rm{K}$ in in-situ conditions. Here, $\sigma_{z}=\sqrt{k_{B}T/m\omega_{z}^{2}}\approx 1.1$ $m$ and $\sigma_{r}=\sqrt{k_{B}T/m\omega_{\rho}^{2}}\approx 0.12$ $m$, with $m$ in this case denoting the mass of sodium atoms, and $\omega_{z}$, $\omega_{\rho}$ the trapping frequency in axial and radial direction, respectively. The arrows in the right-hand plot indicate the direction of the quantization axis relative to the $y$-direction - the direction of the propagation of light - along the center of the trap (y=0,z=0y=0,\ z=0).
@@ III.3 Rotation Matrices
where $B_{\mathrm{tot}}=\sqrt{B_{x}^{2}+B_{y}^{2}+B_{z}^{2}}$. To rotate the field vector on the $y$ axis, a rotation by $\phi$ about $\hat{\mathbf{y}}$ is applied followed by a rotation by $\theta$ about $\hat{\mathbf{z}}$. The combined transformation in Euler‐angle convention [^2] Euler angles $(\alpha,\beta,\gamma)$ define rotations in the Wigner DD‐matrix formalism. Note that $\beta$ above (deviation from the $y$ axis) differs from the Euler angle $\beta$. is
@@
with the upper-left element representing $m^{\prime}=-1$, m=1m=-1. Applying this transformation to the density matrix yields
@@
where the population is initially in the m=1m=-1 state. Equation (20) then becomes:
@@
which satisfy $\sum_{m}\Pi_{m}=1$. For $\theta=0$, the entire population occupies m=1m=-1; for $\theta=\pi$, $m=+1$; and for $\theta=\pi/2$, one finds $\Pi_{\pm 1}=1/4$ and $\Pi_{0}=1/2$.
@@ III.4 Accumulated Phase and Column Densities
Figure 4: Simulated phase delay contours for a) $\sigma_{+}$ and b) $\sigma_{-}$ polarizations showing an apparent displacement of the phase centroid in a thermal atomic cloud of $3\times 10^{8}$ atoms at $30\ \rm{\mu K}$. The dashed white line indicates its Gaussian widths. c) Vertical and d) horizontal line cuts through the center of the clouds highlight an apparent radial shift of about $\pm 40\ \mu\text{m}$ for $\sigma^{\pm}$-polarized light, relative to the center of the ground truth (GT) phase delay for $\beta=\pi$. Note the different x-axis scales in the plots in c) and d) due to the cigar-shaped geometry of the cloud. The vertical cut is scaled for better visualization of the displacement.
@@
with radial and axial widths $\sigma_{r},\sigma_{z}$, the spatial dependence of $\Pi_{m}(\beta)$ induces an asymmetric phase profile for circularly polarized light. Specifically, for right- versus left-handed circular polarization, the phase contour is displaced oppositely along the radial axis. The simulation in Fig. 4 illustrates this effect for a cloud of $3\times 10^{8}$ atoms at $30\,\mu\text{K}$, exhibiting a radial shift of approximately $\pm 40\ \mu\text{m}$ relative to the centroid of the phase distribution at $\beta=\pi$. Although an axial shift is, in principle, also expected, it depends on the magnetic-field gradient along the axial direction, which is negligible near the trap center. As a result, no significant axial displacement is observed in the simulations or in the experiment.
@@ IV Comparison to the Experimental Data
Figure 5: Displacement between the centers of the two circular polarization components as a function of temperature. For each image taken during cooling, the temperature is extracted from fits to the data, and the corresponding displacement is computed using the model described above. At high temperatures, the model predicts the behavior well, but it seems not to fully account for the shift at lower temperatures. The solid black dots show the data; empty dots with a dashed red guide show the simulation.
@@
$\times\Bigl{(}1-\tfrac{1}{2}\sin^{2}\beta(x,y,z)\Bigr{)},$ (29)
@@ V Conclusion
We have introduced a compact three–dimensional framework that links the local tilt of the magnetic field to the phase accumulated by the circular components of an off-axis probe. By combining tensor polarizability with field–frame rotations, the model predicts both phase and spin density transformations, provided $\mathbf{B}(x,y,z)$ is known. The theoretical predictions follow experimental SOAH data down to a few $\mu K$, and reveal a systematic displacement of roughly $\sim 1\,\mu\text{m}$ that persists even in a uniform dipole trap. For small clouds and for $\beta\ll 1$, this factor can be seen as roughly linear, allowing for a fast correction even when the magnetic fields are not precisely known. When the magnetic field is known exactly, this method allows for higher-order aberration corrections down to the diffraction limit.